Find the domain and range of the quadratic equation f(x)=x^2-8x+9.
Answer by stanbon(57361)

Find the domain and range of the quadratic equation f(x)=x^2-8x+9.
----------------
Domain: All Real Numbers
Range:
The minimum point is at (-b/2a, f(-b/2a)
-b/2a= 8/2 = 4
f(4) = 16-21+9 = 4
So Range is "all Real Numbers greater than or equal to 4".
====================
Cheers,
Stan H.